OpenStudy (anonymous):

one more question :) what would be the first step in solving [(x+5)^1/2]-2=(x-1)^1/2

OpenStudy (anonymous):

would do the square root?

OpenStudy (anonymous):

square both sides is the first step

OpenStudy (anonymous):

does that drop the exponent, and you are left with what is inside the parenthesis?

OpenStudy (anonymous):

\[\sqrt{x+5}-2=\sqrt{x-1}\]

OpenStudy (anonymous):

square and get \[x+5-4\sqrt{x+5}+4=x-1\]

OpenStudy (anonymous):

or \[x+9-4\sqrt{x+5}=x-1\] then cancel the \(x\) \[9-4\sqrt{x+5}=-1\] subtract \(9\) \[-4\sqrt{x+5}=-10\] divide by \(-4\) and get \[\sqrt{x+5}=\frac{5}{2}\] then square again and solve for \(x\)

OpenStudy (anonymous):

ok then isolate the radical like in the previous question and then square again?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

only need to square one more time now get \(x+5=\frac{25}{4}\) then subtract \(5\)